OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log 1/3 x. Part 1: Describe the domain, range, and general shape of this graph. (4 points) Part 2: Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts.

OpenStudy (anonymous):

\[f(x)=\log\left({1\over 3x}\right) \] is that it?

OpenStudy (anonymous):

A useful thing to know is \[\log(f/g) = \log(f) - \log(g)\]

OpenStudy (anonymous):

can any of you guys help me with this @ZeHanz and @Directrix

OpenStudy (anonymous):

Is the function the same as the one electrokid typed up? @Tati_Lee

OpenStudy (anonymous):

no it would be f(x)= log(x)/log(1/3) @genius12 i think

OpenStudy (anonymous):

so, you have it as \[f(x)=\log\left({1\over3}x\right)\]?

OpenStudy (anonymous):

I am asking that you typed up y = log 1/3 x in your question. Is that function given in the form electrokid typed it up as?

OpenStudy (zehanz):

So you have \(\log_{\frac{1}{3}}x\). The graph is decreasing (base < 1).

OpenStudy (anonymous):

@ZeHanz We aren't sure whether 1/3 is the base or not which is why I'm trying to ask @Tati_Lee

OpenStudy (zehanz):

@genius12 : I understand, but she wrote f(x)= log(x)/log(1/3).

OpenStudy (anonymous):

I know lol which kinda confused but ok. I guess we can assume that.

OpenStudy (anonymous):

isnt that right?

OpenStudy (anonymous):

Yes it is. If 1/3 is that base that is correct.

OpenStudy (anonymous):

it is the base

OpenStudy (anonymous):



OpenStudy (zehanz):

Domain of any log function is all positive real numbers (as long as no translations to the right or left have been applied).

OpenStudy (zehanz):

Here is a graph: